Solve the system using ROW OPERATION. Please show work.
Thanks!:
4x + y = 6
6x + 2y = 8

Answer by stanbon(75887)

Solve the system using ROW OPERATION. Please show work.
Thanks!:
4x + y = 6
6x + 2y = 8
--------
Multiply thru the 1st row by 3/2 to get:
6x + (3/2)y = 9
6x + 2 y = 8
-----
Subtract 1st row from 2nd row to get:
6x + (3/2)y = 9
0 + (1/2)y = -1
------
Multiply thru 2nd equation by 2 to get:
6x + (1/2)y = 9
0 + y = -2
-----
Subtract (1/2) the 2nd equation from the 1st::
6x + 0 = 10
0 + y = -2
----
Divide thru the 1st equation by 6 to get::
x + 0 = 5/3
0 + y = -2
------------------
Cheers,
Stan H.
---------------